FEL Upgrade Project Weekly Brief
December 3-7, 2001



Highlights:
Three members of the FEL team were elected as APS fellows this year.  Fred Dylla,
George Neil, and Geoff Krafft were all honored for contributions related to their work on the IR Demo.
Congratulations!

We received the bids back for the fabrication of the largest magnets in the FEL Upgrade-the recirculation
dipoles (DY). We are pleased to have received several bids within the price range of our estimated costs.

We thank Charlie Reece for overseeing an effort to Helium and RF pulse process cavity 3 in the FEL injector
cryomodule.  A significant improvement in the usual gradient was obtained. See the operations section for
details.

Management:
An internal review was held this week on optical cavity diagnostics.  See the optics section for details.

The project monthly report for the month of October was completed and distributed to the DOE and ONR
contract monitors.  Some extra time was taken to prepare the financial templates for the Phase 2 IR Upgrade
and UVFEL portions of the project, since October is the first full month of cost performance reporting on these elements.

We are pleased to note that an article on FELs will appear in the January 2002 issue of Physics Today, written
by W. Colson, E. Johnson, M. Kelley, and A. Schwettman.  Michael Kelley is with our group and the others are all longtime collaborators.

WBS 3 (Beam Physics):
UV bypass design is approaching completion. A minor rearrangement was required in the first matching
telescope to alleviate geometric aberrations.  This change will allow common layouts for the first and last
telescopes.

Engineering support focused on "making it all fit".  Most of the 10 lbs. of fertilizer appears to be in the 5 lb. sack, but open it with caution.

A magnet nano-review checked GW and octopole parameters.  They look good.

WBS 4 (Injector):
The modified shield assembly has been assembled and is ready for testing. Drawings for the shield operator
(back end ) and the gun stand are ready for sign off. Testing on the implanted electrode has begun and is so far looking good having been conditioned to 20 Mv/m while drawing an average of 15 pA.

WBS 6 (RF):
Injector HVPS Tank - No known progress.

Quarter Klystrons - Tried to recover 50 kW klystron S/N 007 before upgrading HVPS to no avail.  It appears to arc internally from 7 to 20 kV.  It will need rebuilding by others to make it operational again.  The LCW piping on the first 110 kW cart was tested this week.  New pipe joints are being made today.  The LCW flow rates are adequate to test the 100 kW klystrons at half power.  Larger taps into the building LCW pipes for the final installation.

Quarter HVPS - We are still looking for space and ways to work on the HVPS's.  The circuit design of the HV section is nearly finalized.

WBS 8 (Instrumentation):
The VME Interface card design is complete and ready for documentation.

A camera  has been mounted in the vault and connected to the web for anyone who wishes to view real-time
video on the upgrade.  A link can be found at: http://laser.jlab.org/video/upgrade.

Design of new stepper motor control chassis is underway.  This will include not only an upgraded version, but
additional channels as well.

The second Picomotor Mux chassis has been tested and is ready for use when it is needed.

Documentation is complete for the new Beamviewer Lamp/Camera Power Supply boxes.  These will be sent
out for bid on assembly next week.

WBS 9 (Transport):
Dipoles
Optical Chicane Dipoles (GW)
o  We inspected the initial winding of the first coil as well as bend up fixtures and the potting fixture
    machining at Magnet Enterprises International last week.  Potting will be observed next week by an
    epoxy expert, Jeffrey Hubrig of Innovation Services.
Injector Dipoles (DU/DV)
o  DULY is almost finished revising small portions of the final assemblies.  The specifications are
    being written to match the drawings and revised with what we have learned from our procurements
    of magnets to date.
Arc Dipoles (GY, GX, GQ)
o  The magnetic model of the Bend and Reverse Bend Dipoles (GX & GQ) is being exercised to is
    being exercised to increase the width of the of the good field region per David Douglasí
    specification. The GX is nearly through and the GQ is starting.  The drawings will be revised when
    the adjustment is over.
o  The bids for the 180° Dipoles (DY) are due on Monday.
Magnet Measurement
o  The parts for the dipole measurement stand are signed off
o  We continued work on the probe that will measure the sextupole and the octupole.
Quadrupoles
3 inch quad (QX)
o  The core vender, New England Techni-Coil is making the final adjustment to the 45 degree chamfer
    on the pole edge of the first article core.
o  We resolved the non problem of turn to turn shorts in the coils by unwrapping the copper tubing of
    the cooling circuits.  The wrapped configuration was giving false indications of shorts.
o  A second potential problem is being resolved by a thermal test.  Sporadic brazing of the tubes to
    cooling plates on the initial series of coils is being compared in a thermal test with coils that have
    continuous brazing.
o  A shipment of 64 fully qualified coils is being readied at Magnet Enterprises International.
Trim Quad (QT)
o  The bids were received and evaluated.  The low bidder is being asked to detail a cost breakdown
    because of the spread in the quotations.
Sextupole (SF)
o  DULY Research continues to model the sextupole.
Octupole (OT)
o  Work on this magnet is on furlough.
Corrector Dipoles (DB, DJ)
o  Milhous Co. is about 6 shy of delivering the entire order DB and DJ coils.
Beam Line and Vacuum
o  Dave Waldman chaired the kick-off meeting to start the task order with AES to design the Arc
    Chambers.
o  Several sample chambers with the thicker plates necessary for the upgrade have been made to test
    out the magnetic properties.
o  Design work on girders and stands continued. All details of the region between the cryomodules are
    now complete.
o  All return leg girders have been received.
o  All the new three leg pedestals are now out for bid.

WBS 10 (Wiggler):
Magnet measurement scans were taken on wiggler 1 to check the corrector values to be used during
commissioning and to better verify the vertical trajectory. Analysis work was done on the affect on gain of using the dispersion section at low field.  This is the regime where analytical models do not apply.  This is quite
important for operation at full power.

WBS 11 (Optics):
Progress for the upgrade was made in the following areas:

The detailing of the design for the optical cavities continues; we are 90% complete on the drawing package for
components from the "big C" (the yaw axis yoke) through the mirror holders.

Testing of the ultraviewer has been delayed due to illness of one of the staff.  We plan to start Friday (12/07)
afternoon.  The OCMMS tests continue.  An internal review of the OCMMS beamline hardware revealed a
possible susceptibility to mirror movement due to power absorbed from the FEL harmonics.  We are redesigning that part.

Discussions with a potential vendor for the drive laser were held.  We received an updated budgetary quote
from a vendor on the ultrafast laser.  Calculations and planning continue on enhanced FEL pulsewidth
diagnostics.

Operations/Decommissioning/Other Activities:
From C.Reece's report of RF conditioning of the injector cryounit:

A note describing the recent test of pulsed rf helium processing on FEL I-3 is available in DocuShare at:
http://docushare.jlab.org/Get/File-4729/Pulsed_rf_helium_.pdf

The cavity showed significant improvement with cw processing and also subsequently with 2 ms pulsed
processing. In pulsed mode the cavity was limited by prompt arc trips at 17.5 MV/m.  In cw mode, the cavity
was limited to 15 MV/m by IR trips. It was initially limited by field-emission induced quench at 12 MV/m.
Operationally, it has been limited to ~9.2 MV/m by field emission loading. It now appears quite usable to at least 13 MV/m.

Thanks to J. Gubeli's efforts, all FEL optics associated with the optical cavities have been removed and are
either stored in User Lab 2 or are awaiting survey by RadCon.  Before disassembly, a videotape with narration
was made.  All of the turning mirrors were removed, their orientation marked, and their flatness measured with
the Wyko laser interferometer for later analysis.  We removed, disassembled, and stored for later use the optical transport system beamtubes.  S. Benson removed the lightbox yoke and stored the associated components.  The piece of drive laser optical transport that places the Brewster window close to the light box was removed.

We gave a tour to staff from NSWC Dahlgren.